OpenStudy (josee):

Find the slope of the line tangent to the graph of ln(xy)-2x=0 when x= -1

OpenStudy (anonymous):

(y+xy')/xy -2 =0 ===> y' =-1 ===> y'(x=-1) =-1

OpenStudy (josee):

um..these are the answer choices.. 1. slope = 3/2e^-2 2. slope= 3/2e^2 3. slope= 3e^-2 4. slope= -3e^2 5. slope= -3/2e^2 6. slope= -3e^-2

OpenStudy (josee):

lol :P